# Symmetry and Topology in the Non-Hermitian Kitaev chain

### v2

@@ IV Conclusion
With recent experimental realization of non-Hermitian skin effect in mechanical metamaterials [22, 63], topoelectric circuits [81, 18] and photonic systems [64, 67, 80, 76, 21, 79, 62], the interplay between topology and non-Hermiticity has become experimentally relevant. The versatility of these platforms allows for the engineering of a wide range of non-reciprocal couplings and effective interactions [78], making the realization of non-Hermitian Majorana physics a tangible prospect. In this work, we have provided a comprehensive characterization of the non-Hermitian Kitaev chain over the full complex parameter space, establishing a consistent correspondence between bulk topology, zero-energy Majorana modes, and phase boundaries. For specific parameter regimes, we obtained analytical expressions for points on the topological phase boundary, as illustrated in Fig. 5, offering concrete insight into the structure of the phase diagram. It remains to be seen whether the phase diagram can be determined analytically in full generality, including a characterization of the shape of the phase boundary; nevertheless, our results lay a systematic and broadly applicable foundation for exploring non-Hermitian topological phase transitions in this and related models.
@@ Acknowledgements
Appendix A Symmetry analysis for non-Hermitian Kitaev chain
@@
$H^{\text{NH}}$ $=-\mu\sum_{j=1}^{L}c_{j}^{\dagger}c_{j}-\frac{1}{2}\sum_{j=1}^{L}(t_{1}c_{j}^{\dagger}c_{j+1}+t_{2}c_{j+1}^{\dagger}c_{j}+\Delta_{1}e^{-i\varphi_{1}}c_{j}c_{j+1}+\Delta_{2}e^{i\varphi_{2}}c_{j+1}^{\dagger}c_{j}^{\dagger})\,$, (A1)
@@
$H(k)=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (A2)
and $\mathcal{C}_{k}^{\dagger}=\begin{pmatrix}c_{k}^{\dagger}&c_{-k}\end{pmatrix}$. The discrete symmetries have been discussed in [33]. The spectral constraints arising from the possible symmetries of the system are summarized in the Table A1.
Table A1: Symmetries and their implications on the spectrum in momentum space
@@
A.1 Time reversal symmetry (TRS)
@@
$H(-k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{-ik}+t_{2}e^{ik}+2\mu\end{array}\right)$ (A6)
$H(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (A9)
@@
$H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}\end{array}\right)$ (A12)
$H^{\dagger}(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu^{*}\end{array}\right)$ (A15)
@@
$\quad+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (A18)
@@
$\quad+4t_{1}^{*}t_{2}^{*}+8\Delta_{1}\Delta_{2}e^{i\varphi_{1}-i\varphi_{2}}\sin^{2}k$ (A19)
@@
$\quad+4\Delta_{2}^{2}\sin^{2}k+4\Delta_{1}^{2}\sin^{2}k+8|\mu|^{2}$ (A20)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A21)
@@
A.2 Particle hole symmetry (PHS)
@@
$-H(-k)$ $=\left(\begin{array}[]{cc}t_{1}e^{ik}+t_{2}e^{-ik}+2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu\end{array}\right)$ (A25)
$H(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (A28)
@@
$-H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}\end{array}\right)$ (A31)
$H^{\dagger}(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (A34)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (A37)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (A38)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A39)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A40)
@@
A.3 Adjoint time reversal symmetry (TRS)
@@
$H(-k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{-ik}+t_{2}e^{ik}+2\mu\end{array}\right)$ (A43)
$H(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (A46)
@@
$H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}\end{array}\right)$ (A49)
$H^{\dagger}(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (A52)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (A55)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (A56)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A57)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A58)
@@
A.4 Adjoint particle hole symmetry (PHS)

### v3

@@ IV Conclusion
The interplay of topology and non-Hermiticity is very interesting and gives rise to many new theoretical ideas such as the non-Hermitian skin effect and non-Bloch bulk-boundary correspondence. There are also numerous experimental validations of these ideas in mechanical metamaterials [23, 67, 57], topoelectric circuits [85, 19], and photonic systems [68, 71, 84, 80, 22, 83, 66]. The key idea in all such experimental platforms is not to deal with real quantum Hamiltonians, but only to mimic the physics of a quantum Hamiltonian. This is achieved by tuning various electrical components such as resistors, inductors and capacitors in topoelectric circuits [85], by artificially engineering non-reciprocal hopping amplitudes in active mechanical circuits [23], or by modulating gain and loss in unidirectional coupling link rings in optical systems [84]. Additionally, non-Hermitian physics has also found application in designing new highly sensitive sensors, as demonstrated in Ref. [46]. The versatility of these platforms allows for the engineering of a wide range of non-reciprocal couplings and effective interactions [82]. Realizing a non-Hermitian Kitaev chain may be challenging because of the superconducting pairing terms. However, with all the advances in the experimental field, it is not unreasonable to be optimistic that such interactions may be realized, or at least the underlying physics may be mimicked in an experimental platform.
Overall, in this work we have provided a comprehensive characterization of the non-Hermitian Kitaev chain over the full complex parameter space, establishing a consistent correspondence between bulk topology, zero-energy Majorana modes, and phase boundaries. For specific parameter regimes, we obtained analytical expressions for points on the topological phase boundary, as illustrated in Fig. 5, offering concrete insight into the structure of the phase diagram. It remains to be seen whether the phase diagram can be determined analytically in full generality, including a characterization of the shape of the phase boundary; nevertheless, our results lay a systematic and broadly applicable foundation for exploring non-Hermitian topological phase transitions in this and related models.
@@ Acknowledgements
Appendix A Probability densities for ground and excited states
In this appendix, we elaborate on obtaining the probability densities of the zero energy modes ($E=0$) and the first excited states (with energy $\pm E$). By first excited state we mean eigenstates with lowest magnitude of energy with $|E|\neq 0$. Consider the Hamiltonian discussed in the main text, Eq. (A1) as follows,
$H^{\text{NH}}=\frac{1}{2}$ $\Psi^{\dagger}\begin{pmatrix}A&B&0&0&0&\cdots\\ C&A&B&0&0&\cdots\\ 0&C&A&B&0&\cdots\\ 0&0&C&A&B&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\\ \end{pmatrix}\Psi,$ (A1)
where $\Psi^{\dagger}=\begin{pmatrix}c_{1}^{\dagger}&c_{1}&c_{2}^{\dagger}&c_{2}&\cdots\end{pmatrix}$ and $A$, $B$ and $C$ are $2\times 2$ matrices,
$A=\begin{pmatrix}-\mu&0\\ 0&\mu\end{pmatrix}\qquad B=\begin{pmatrix}-\frac{t_{1}}{2}&\frac{\Delta_{2}}{2}\\ -\frac{\Delta_{1}}{2}&\frac{t_{2}}{2}\end{pmatrix}\qquad C=\begin{pmatrix}-\frac{t_{2}}{2}&-\frac{\Delta_{2}}{2}\\ \frac{\Delta_{1}}{2}&\frac{t_{1}}{2}\end{pmatrix}$,
Let us start with an eigenvector of the Hamiltonian corresponding to energy $E$. Denote the $(2j-1)^{\text{th}}$ component of the eigenvector as $u_{j}$ and $2j^{\text{th}}$ component as $v_{j}$ where $j\in{1,2,\cdots,L}$. We identify the odd component of the eigenstate $\psi_{1,j}=u_{j}c_{j}$ to be the corresponding component for the particle wavefunction and similarly $\psi_{2,j}=v_{j}c_{j}^{\dagger}$ is the component of the hole wavefunction. Thus, the probability density of the particle with energy $E$ at $j^{\text{th}}$ site is given by $|u_{j}|^{2}$ and for the hole, density at $j^{\text{th}}$ site is $|v_{j}|^{2}$.
For the eigenstates of the Hamiltonian corresponding to the energy eigenstates $\pm E$, let us denote the particle and hole probability density at $j^{\text{th}}$ site with $|u^{\pm}(j)|^{2}$ and $|v^{\pm}(j)|^{2}$. The particle-hole symmetry (PHS) of the non-Hermitian Kitaev chain implies that if a particle with energy $\pm E$ localizes at one end of the chain, then a hole of energy $\mp E$ localizes at the other end (see Sec. III.3). Now, the probability densities of these two wavefunctions $\psi_{\pm}$ at site $j$ are given by
$|\psi_{\pm}(j)|^{2}=|u^{\pm}_{j}|^{2}+|v^{\pm}_{j}|^{2}.$ (A2)
As a result of PHS, probability density of particles $|u^{+}|^{2}$ with energy $E$, has an equal and opposite localization as compared to the probability density of holes $|v^{-}|^{2}$. Equivalently, $|u^{-}|^{2}$ has an equal and opposite localization as compared to $|v^{+}|^{2}$. This immediately leads to the conclusion that the probability density $|\psi^{+}|^{2}$ of the wavefunction corresponding to energy +E+E has an equal and opposite localization as compared to the probability density $|\psi^{-}|^{2}$ of the wavefunction corresponding to energy E-E cf. Fig. A1(b).
In the case of Majorana modes, we present the combined probability density i.e., summing the probability density for both the zero modes as follows,
$|\psi_{0}^{(1)}|^{2}+|\psi_{0}^{(2)}|^{2}=|u_{0}^{(1)}|^{2}+|v_{0}^{(1)}|^{2}+|u_{0}^{(2)}|^{2}+|v_{0}^{(2)}|^{2},$ (A3)
where $u_{0}^{(1,2)}$ and $v_{0}^{(1,2)}$ are the particle and hole components corresponding to the two Majorana zero modes $\psi_{0}^{(1,2)}$. It is now evident that the symmetric profile of the combined probability density of the Majorana zero modes is also a consequence of PHS cf. Fig. A1(a).
Figure A1: Probability density of zero modes and excited states of the non-Hermitian Kitaev chain with $L=50$ sites under open boundary conditions and $\mu=0.2+i,t_{1}=1+i,t_{2}=1.5+2i,\Delta_{1}=2+i,\Delta_{2}=1+3i$. (a) The total probability density of the two Majorana zero modes $\psi_{0}^{(1)}$ and $\psi_{0}^{(2)}$ is symmetric and localized at the edges. (b) $\psi_{1}^{+}$ and $\psi_{1}^{-}$ are the energy eigenstates corresponding to complex energies $\pm E,\text{Re}(E)>0$, where $|E|$ is the smallest absolute value of energy in the spectrum. They see an equal and opposite skin effect localization.
Appendix B Symmetry analysis for non-Hermitian Kitaev chain
@@
$H^{\text{NH}}$ $=-\mu\sum_{j=1}^{L}c_{j}^{\dagger}c_{j}-\frac{1}{2}\sum_{j=1}^{L}(t_{1}c_{j}^{\dagger}c_{j+1}+t_{2}c_{j+1}^{\dagger}c_{j}+\Delta_{1}e^{-i\varphi_{1}}c_{j}c_{j+1}+\Delta_{2}e^{i\varphi_{2}}c_{j+1}^{\dagger}c_{j}^{\dagger})\,$, (B1)
@@
$H(k)=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (B2)
and $\mathcal{C}_{k}^{\dagger}=\begin{pmatrix}c_{k}^{\dagger}&c_{-k}\end{pmatrix}$. The discrete symmetries have been discussed in [34]. The spectral constraints arising from the possible symmetries of the system are summarized in the Table B1.
Table B1: Symmetries and their implications on the spectrum in momentum space
@@
B.1 Time reversal symmetry (TRS)
@@
$H(-k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{-ik}+t_{2}e^{ik}+2\mu\end{array}\right)$ (B6)
$H(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (B9)
@@
$H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}\end{array}\right)$ (B12)
$H^{\dagger}(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu^{*}\end{array}\right)$ (B15)
@@
$\quad+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (B18)
@@
$\quad+4t_{1}^{*}t_{2}^{*}+8\Delta_{1}\Delta_{2}e^{i\varphi_{1}-i\varphi_{2}}\sin^{2}k$ (B19)
@@
$\quad+4\Delta_{2}^{2}\sin^{2}k+4\Delta_{1}^{2}\sin^{2}k+8|\mu|^{2}$ (B20)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B21)
@@
B.2 Particle hole symmetry (PHS)
@@
$-H(-k)$ $=\left(\begin{array}[]{cc}t_{1}e^{ik}+t_{2}e^{-ik}+2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu\end{array}\right)$ (B25)
$H(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (B28)
@@
$-H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}\end{array}\right)$ (B31)
$H^{\dagger}(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (B34)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (B37)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (B38)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B39)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B40)
@@
The generator of PHS for the non-Hermitian Kitaev chain is $\sigma_{x}$, i.e. we have,
$\sigma_{x}H(k)^{T}\sigma_{x}=-H(-k).$ (B41)
In the most general parameter regime, PHS is the only symmetry of the non-Hermitian Kitaev chain, Table 1. Following the symmetry-based topological classification of [35], we observe that the generator of PHS squares to identity, i.e. $\sigma_{x}^{2}=\mathbbm{1}$. Using this and the fact that the spectrum has a line gap, we conclude that the non-Hermitian Kitaev chain lies in class D, which has a $\mathbb{Z}_{2}$ topological invariant.
B.3 Adjoint time reversal symmetry (TRS)
@@
$H(-k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{-ik}+t_{2}e^{ik}+2\mu\end{array}\right)$ (B44)
$H(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (B47)
@@
$H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}\end{array}\right)$ (B50)
$H^{\dagger}(k)^{T}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (B53)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (B56)
@@
$+8\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (B57)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B58)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B59)
@@
B.4 Adjoint particle hole symmetry (PHS)